# How Much Content Do LLMs Generate That Induces Cognitive Bias in Users?

### v1

@@ 9 Limitations
This study examines the extent to which LLM outputs alter original content in ways that may influence human judgment. While we ground our assumptions in prior cognitive science research, as discussed in the paper, we aim to validate these effects through a dedicated user study measuring how such alterations impact human decision-making.
Most of our mitigation methods were designed with one specific issue in mind (e.g. framing, hallucinations, or primacy effects). In our future work, we aim to explore more general mitigation techniques that address overlapping biases simultaneously.
Our work aims to encourage the NLP community to continuously test LLMs for biases that can influence users’ decision-making, which is why we introduce the self-updating NewsLensSync dataset Akshaya (2025). Currently the dataset holds articles related to "political" domain. Soon we will try to incorporate other areas to make it more generalizable and usable across diverse set of researchers. However, our data could be misused for training models to reinforce existing ideological narratives or use falsified data. We urge researchers to be cautious in framing and contextualizing their use of the data. We provide our data under the CC-BY-4.0 license.
@@
Our analysis shows that LLMs pay more attention to content from the beginning of the context, hallucinate facts, and alter the text framing in their outputs. These alterations can have an impact on the user, and skew their decision-making. These findings underscore the risks of working with LLMs, even in everyday tasks.
Experiments. All experiments are run on NVIDIA RTX A6000 for open-source models and the official APIs for closed-source models with a fixed random seed. Our run time was approximately 360 GPU hours. All prompts details are included in Appendix E.
Appendix A LLMs as a framing judge
Previous work has shown that high-capability LLMs like GPT-4 can achieve over 80% agreement with human experts on question answering tasks, the level of agreement observed between humans themselves (81%). GPT-4 reached 85% agreement with human judgments, suggesting that LLMs can serve as effective proxies for human preferences Zheng et al. (2023b). Thus, we test three high-capability models to classify the framing ( GPT-3.5-turbo, GPT-4-turbo, and GPT-4o-mini). We experimented on a sample of 500 randomly selected Amazon reviews, using the user-provided rating as ground truth to evaluate the accuracy of the models. We asked the models to rate the product based on the review, then we mapped ratings 1 and 2 as negative, 3 as neutral, and 4 and 5 as positive. The GPT-3.5-turbo, GPT-4-turbo, and GPT-4o-mini achieved accuracy of 0.89, 0.91, and 0.92, respectively. Based on this, we selected GPT-4o-mini to serve as the framing judge.
Appendix B Model setup and Hyperparameters
Our implementation utilizes the Hugging Face Transformers library using the official APIs. We employ Flash Attention 2 when available to enhance computational efficiency during inference. For consistency across experimental conditions, we maintain fixed generation parameters with temperature=0.01, do_sample=False, and max_new_tokens= 500 set to a constant with fixed random seed. Table 7 shows the models we evaluated, their context window, and the rationale behind choosing them.
@@ Appendix B Model setup and Hyperparameters
Table 7: Models used in our evaluation, along with their context-window sizes and selection rationale.
Appendix C Detailed Description of Mitigation Methods
Table 8 provides a concise, side-by-side reference for all 18 mitigation approaches evaluated in this work, including their intended bias target(s), operational definition, key hyper-parameters, and the original citation.
Family Method Operational definition Target bias Key hyper-parameters Citation
@@ Appendix C Detailed Description of Mitigation Methods
Epistemic Tagging Requires models to express confidence levels alongside factual assertions, creating a more nuanced representation of the model’s knowledge state. Forces the model to evaluate its own fact-checking abilities before responding and provides users with meta-cognitive signals about response reliability. Hallucination Confidence level scales - in this case we used High and Low. This can incude moderate, very high, very low depending on task. Lin et al. (2022)
@@
Weighted Summaries We pre-allocate the summary length budget so that 33% focuses on the beginning, 34% on the middle, and 33% on the end. The model is instructed not to exceed these chunk limits, which helps prevent overemphasis on earlier parts. Primacy Token ratio $0.33{:}0.34{:}0.33$. inspired by Liu and Lapata (2019)
Re-rank Attention-Sort Re-ordering Run a forward pass (without generating the final summary) to estimate cross-attention weights for each paragraph or segment, then reorder segments from lowest-attended to highest-attended. The model is then re-prompted with this new order, which pushes under-attended parts later in the context and encourages more balanced coverage. Primacy 2 iterations; paragraph granularity.
@@
Rejection Sampling If top-1 token would increase chunk imbalance, set its logit to $\infty$- ∞ and resample within top-$k$. Primacy $k=5$.
Self-Debias Decoding Dual-pass decoding: at each step run a second forward pass on the current context preceded by a bias-inducing prefix that names the undesired attribute; obtain $p_{\text{bias}}$, compute $\Delta=p_{\text{main}}-p_{\text{bias}}$, and down-scale tokens with $\Delta<0$ via $\alpha(\Delta)=e^{\lambda\Delta}$. Framing $\lambda=10$; bias prefix $<\!30$ tokens; refresh every 4 steps. Schick et al. (2021)
@@
Table 8: Overview of the 18 bias-mitigation strategies evaluated in this work. Methods are grouped by the family that they belong to, annotated with the bias they aim to address, and accompanied by the main hyper-parameters and primary citation(s) used in our experiments.
Appendix D Full Coverage, Framing, and Bias Metrics
Table 9 reports the complete results for all models and mitigation strategies.
@@ Appendix D Full Coverage, Framing, and Bias Metrics
Table 9: Coverage, framing-change, and primacy-bias metrics for Amazon Reviews and MediaSum News. Metrics: $\varphi_{\mathrm{frame}}$ (framing-change fraction, $\downarrow$), $\bar{s}_{b},\bar{s}_{m},\bar{s}_{e}$ (cosine similarity with the first/middle/last third of the source, $\uparrow$), $\psi_{\mathrm{pri}}$ (share of summaries whose similarity to the beginning exceeds that to the middle by $>5\%$, $\downarrow$), and $\rho_{\text{pri}}$ (share of summaries exhibiting primacy bias, $\downarrow$).
Appendix E Prompt Templates Used in Our Experiments
Table 10 lists the exact prompt instruction templates supplied to each model under all the experimental settings.
@@ Appendix E Prompt Templates Used in Our Experiments
Self-Awareness Prompting “You are an unbiased summarizer. Be mindful not to introduce any framing bias or omit the middle. Preserve the original sentiment.
Please summarize the following text: [DOCUMENT_TEXT]
@@
[PARTIAL_SUMMARY_1]
[PARTIAL_SUMMARY_2]
@@
Table 10: Prompt strategies and the prompts used in our experiments.

### v2

@@ 10 Limitations
This study investigates the degree to which LLM outputs modify original content in ways that may impact human judgment. Most of our mitigation methods were designed with one specific issue in mind (e.g., framing, hallucinations, or primacy effects). In our future work, we intend to explore more general mitigation techniques that address overlapping biases simultaneously.
Our current primacy bias score uses a fixed, generic threshold to identify summaries that disproportionately focus on the beginning of the source text. While this threshold provides a consistent way to detect bias, our future work aims to explore ways to refine it. In addition, our evaluation relies on one LLM to assess another LLM’s outputs in the framing bias metric. Although commonly used, LLM-as-a-judge approaches have drawbacks. We use LLM-as-a-judge to the best of our knowledge, given current best practices.
Our work aims to encourage the NLP community to continuously test LLMs for biases that can influence users’ decision-making, which is why we introduce the self-updating NewsLensSync dataset. We urge researchers to exercise caution when framing and contextualizing their use of the data. We provide our data under the CC-BY-4.0 license.
@@
Our analysis reveals that LLMs tend to focus more on content from the beginning of the context, exhibit hallucination of facts, and alter the text framing in their outputs. These alterations can affect the user and influence their decision-making. These findings highlight the risks associated with working with LLMs, even in everyday tasks.
Experiments. All experiments are run on NVIDIA RTX A6000 for open-source models and the official APIs for closed-source models with a fixed random seed. Our run time was approximately 360 GPU hours. All prompts details are included in Appendix A.8.
Appendix A Appendix
A.1 LLM as a framing judge
Previous work has shown that high-capability LLMs like GPT-4 can achieve over 80% agreement with human experts on question-answering tasks, with the level of agreement observed among humans themselves being 81% cite zheng2023judging. GPT-4 reached 85% agreement with human judgments, suggesting that LLMs can serve as effective proxies for human preferences Zheng et al. (2023b). Thus, we test three high-capability models to classify the framing in our study ( GPT-3.5-turbo, GPT-4-turbo, and GPT-o4-mini). On a sample of 500 randomly selected Amazon reviews, we use the user-provided ratings as ground truth to evaluate the models’ accuracy. We ask the models to rate the product based on the review, then we map ratings 1 and 2 as negative, 3 as neutral, and 4 and 5 as positive. The GPT-3.5-turbo, GPT-4-turbo, and GPT-o4-mini achieve accuracy of 0.89, 0.91, and 0.92, respectively. Based on this, we select GPT-o4-mini to serve as the framing judge. We prompt the model as follows:
“[TEXT]
Classify the framing of the text as Positive, Negative, or Neutral. Respond with the class label only.”
to obtain the text framing.
A.2 Model Setup and Hyperparameters
Our implementation uses the Hugging Face Transformers library via its official APIs. We employ Flash Attention 2 when available to enhance computational efficiency during inference. For consistency across experimental conditions, we maintain fixed generation parameters with temperature=0.01, do_sample=False, and max_new_tokens= 500 set to a constant with a fixed random seed. Table A1 presents the models we evaluate, their context windows, and the rationale for choosing them.
All aggregate statistics reported in the abstract are computed using a mean-of-means approach across all six models (GPT-3.5-turbo, Llama-3-8B-Instruct, Llama-3.2-3B-Instruct, Phi-3-mini-4k-Instruct, Qwen3-4B-Instruct, Gemma-3-27B-IT) and both datasets (Amazon Reviews and MediaSum News), where each model-dataset pair contributes one mean calculated from 1000 examples.
@@ A.2 Model Setup and Hyperparameters
Qwen3-4B-Instruct 256k A small, high-performing open-source model included to broaden architectural diversity Yang et al. (2025).
Gemma-3-27B-IT 128k A large, state-of-the-art open-source model included to test for scale effects Team et al. (2025).
Table A1: Models used in our evaluation, along with their context-window sizes and selection rationale.
Table A2 lists the cut-off for the models used for our hallucination study.
Model Cut-off date (training data freshness)
Llama3-8b-instruct Mar 2023
Llama 3.2-3b-instruct Oct 2024
Phi-3-mini-4k-instruct Oct 2023
Table A2: Models and their cut-off dates for training data freshness.
A.3 Further Information on our NewLensSync Dataset
We use a fixed list of trusted domains: bbc.co.uk, reuters.com, apnews.com, npr.org, pbs.org, theguardian.com, wsj.com, nytimes.com, and propublica.org.
The NewsLensSync orchestrator builds a datastore for each article containing source, author, title, api_description (from NewsAPI), webscraped_description (extracted directly from the article to overcome the API limit), falsified_description (negated version for hallucination testing), api_content, url, and publishedAt.
A scheduled CRON job performs one API call per day over the trusted list, minimizing duplication. Any rare overlaps are removed during preprocessing, retaining only unique records for LLM evaluation. Duplication checking is deferred to evaluation time to avoid overhead. The dataset continues to self-update for future research use.
A.4 Descriptive Statisics on Amazon Reviews and Information on MediaSum News datasets
Dataset Mean Median SD Min Max
Amazon Reviews 2180.3 2185.5 354.5 790 3169
MediaSum News 1166.1 1054.0 621.8 89 2924
Table A3: Descriptive statistics of text lengths (tokens) for Amazon Reviews and MediaSum News.
A.5 Statistical Test Results for Primacy Bias
Model Dataset T-Statistic P-Value
Llama-3-8B-Instruct Amazon Reviews 28.2896 8.62e-130
Llama-3-8B-Instruct MediaSum News 26.8965 2.64e-120
Llama-3.2-3B-Instruct Amazon Reviews 28.3100 6.26e-130
Llama-3.2-3B-Instruct MediaSum News 25.1535 1.50e-108
Phi-3-mini-4k-Instruct Amazon Reviews 18.2547 1.81e-64
Phi-3-mini-4k-Instruct MediaSum News 19.9294 1.14e-74
Qwen3-4B-Instruct Amazon Reviews 24.9284 4.82e-107
Qwen3-4B-Instruct MediaSum News 21.6017 3.18e-85
Gemma-3-27B-IT Amazon Reviews 27.7788 2.64e-126
Gemma-3-27B-IT MediaSum News 24.1413 1.01e-101
Table A4: Paired t-test results comparing the cosine similarity between the summary and the beginning of the source text ($s_{b}$) versus the middle of the source text ($s_{m}$). The statistics shown are for the baseline performance of each model.
To validate that the observed primacy effect (where summary-to-beginning similarity, $s_{b}$, is higher than summary-to-middle similarity, $s_{m}$) is a systematic bias, we perform paired t-tests for each model and dataset combination. Table A4 presents the detailed statistical results. $p\ll.001$ across all tested models confirms that the higher similarity scores for the beginning segment are statistically significant, and the large t-statistics further indicate a substantial effect size.
A.6 Detailed Descriptions of Mitigation Methods
Family Method Operational definition Target bias Hyper-parameters Citation
@@ A.6 Detailed Descriptions of Mitigation Methods
Epistemic Tagging Requires models to express confidence levels alongside factual assertions, creating a more nuanced representation of the model’s knowledge state. Forces the model to evaluate its own fact-checking abilities before responding and provides users with meta-cognitive signals about response reliability. Hallucination Confidence level scales (we used high and low). This can include moderate, very high, very low depending on task. Lin et al. (2022)
@@
Weighted Summaries We pre-allocate the summary length budget so that 33% focuses on the beginning, 34% on the middle, and 33% on the end. The model is instructed not to exceed these chunk limits, which helps prevent overemphasis on earlier parts. Primacy Token ratio 0.33:0.34:0.330.33{:}0.34{:}0.33. inspired by Liu and Lapata (2019)
Re-rank Attention-Sort Re-ordering Run a forward pass to estimate cross-attention weights for each paragraph or segment, then reorder segments from lowest-attended to highest-attended. The model is then re-prompted with this new order, which pushes under-attended parts later in the context and encourages more balanced coverage. Primacy 2 iterations; paragraph granularity.
@@
Rejection Sampling If top-1 token would increase chunk imbalance, set its logit to -$\infty$ and resample within top-$k$. Primacy $k=5$.
Self-Debias Decoding Dual-pass decoding: at each step run a second forward pass on the current context preceded by a bias-inducing prefix that names the undesired attribute; obtain $p_{\text{bias}}$, compute $\Delta=p_{\text{main}}-p_{\text{bias}}$, and down-scale tokens with $\Delta<0$ via $\alpha(\Delta)=e^{\lambda\Delta}$. Framing $\lambda=10$; bias prefix <30<\!30 tokens; refresh every 4 steps. Schick et al. (2021)
@@
Table A5: Overview of the 18 bias-mitigation strategies in this work. Methods are grouped family, annotated with the bias they aim to address, and accompanied by main hyper-parameters and primary citation(s).
A.7 Full Mitigation Results
@@
Table A6: Coverage, framing-change, and primacy-bias metrics for Amazon Reviews and MediaSum on $\varphi_{\mathrm{frame}}$ (framing-change fraction, $\downarrow$), $\bar{s}_{b},\bar{s}_{m},\bar{s}_{e}$ (cosine similarity with the first/middle/last third of the source, $\uparrow$), $\psi_{\mathrm{pri}}$ (share of summaries whose similarity to the beginning exceeds that to the middle by $5\%$, $\downarrow$), and $\rho_{\text{pri}}$ (share of summaries exhibiting primacy bias, $\downarrow$).
A.8 Prompt Templates
@@
Self-Awareness Prompting “You are an unbiased summarizer. Be mindful not to introduce any framing bias or omit the middle. Preserve the original sentiment. Please summarize the following text: [DOCUMENT_TEXT]
@@
[PARTIAL_SUMMARY_1] [PARTIAL_SUMMARY_2]
@@
Table A7: Prompt strategies and the prompts used in our experiments.
A.9 Experimental Details on Products for Human Study
Table A8: Product reviews and their summaries used for the human-subject experiment.
Product Review Summary
Refrigerator water filter We installed this filter on our Whirlpool Refrigerator and two months later the water was only trickling. I reseated it just in case something had gotten in the line. Nothing had gotten in. The filter was plugged after two months. The valve and all are fine. After reseating water was still trickling. We ordered and installed the factory Whirlpool filter and as expected the water flow is normal. The waster valve made noise in the days that this cheap filter was clogged, think of the damage that does over time! Why take a chance with this cheaper unit coming apart inside and risking a Service Call because it damaged the valve that is hard to get to?????<br /><br />PAY TYHE MONEY. YOU BOUGHT A NICE FRIDGE. BUY THE WHIRLPOOL FACTORY FILTER. IT IS YOUR HEALTH, IF THIS FILTER CLOGGS AFTER TWO MONTHS THEN IT IS FAILING INSIDE. The old whirlpool filter went over one year not changed and never had an issue. I now have the Whirlpool filter in my subscribe-n-save every six months. I initially tried a cheaper filter for my Whirlpool Refrigerator, but after two months it clogged and caused water flow issues. I switched to the factory Whirlpool filter and the water flow returned to normal. It’s worth investing in the quality filter to avoid potential damage and service calls.
Coffee water filter Three stars ONLY because the filter assembly was almost impossible for me to get open. If you can’t replace the filter, the filter assembly is useless. The bottom of the filter assembly requires the sides of the hard acrylic be squeezed together in order for the bottom to be released (refer to the attached photo). After several failed attempts, I was able to get the bottom released by placing it against a hard object and pressing down on the opposite side with both thumbs. Very worrisome because the acrylic could just as easily cracked, ruining the filter assembly. So, the bottom line is if you don’t have strong hands, it will be very hard to open the filter assembly to replace the filter every two months. Otherwise, this starter kit fits nicely in my Keurig 2.0, K300, the water is clean and tastes very good. The reusable K-cup is another plus for the starter kit. My Keurig recognizes it and allows for brewing loose leaf tea or non-K-cup coffee of my choice. The filter assembly was difficult to open, requiring strong hands to avoid cracking the acrylic. However, the water tastes great and the reusable K-cup is a nice addition for brewing loose leaf tea or non-K-cup coffee. Fits well in my Keurig 2.0, K300.
Electric kettle Cons:<br /> * The first thing you will notice on unpacking is the short cord. It is the shortest appliance cord I have ever seen and is quite inconvenient.<br /> * The build of the kettle is not great, especially considering the price. I am worried that the gooseneck will fall off.<br /> * The electronic controls are clumsy when compared to the Cuisinart Perfect Temp simple presets.<br /> * The heating element is small so heating is slower than expected.<br /> * Bonavita designs never get it quite right. I use their coffeemaker which works great but has a smoked glass water reservoir that makes it hard to get an exact water level. The waterproof cover for the electronics on this kettle are frosted instead of clear so they are not easy to see.<br /><br />Pros:<br /> * Heats water to very close to the desired temps.<br /> * Gooseneck is great for pourover coffee preparation.<br /> * Takes up little counter space but must be very close to an outlet. The short cord and build quality are disappointing, but the Bonavita kettle heats water well and is compact. The gooseneck is perfect for pour-over coffee. Controls are not as user-friendly as other brands. Overall, it’s a decent option for those who prioritize temperature accuracy and pour control.
Screen protector Not overwhelmed nor am I disappointed. This protector is on par with many on the market. Installation is all about preparation. The included guide frame is nice and takes guesswork out of alignment. If you’ve prepped the screen and manage to not stir up dust then you should get an installation that only needs bubbles squeezed out. Take your time and you can get a perfect installation. Take your time.<br /><br />That being said, the frame is all that really sets this one apart from others for this iPad and the others we own. You will get the same quality as if you pick another brand as they’re all pretty much the same one in different packaging.<br /><br />It works and looks as it should. After a bit of use you’ll see fingerprints start to show up. It will happen. You can minimize this by using a clean soft cloth with no cleaner. Eventually though it will start to show every print. Up to you to live with it or install the second one. This screen protector is decent, with easy installation using the guide frame. Quality is similar to other brands. Fingerprints show up after use, but can be minimized with a clean cloth. Overall, it works as expected but may require a second installation for a flawless look.
Dryer vent hood The Good:<br />The construction is excellent. Thick stainless steel and well made.<br />The damper works well. We are using this as a dryer vent and the damper moves out fully with even the mildest air flow. Its great not to have a screen that clogs with lint since this is installed on the second floor and getting to the old vent was a big hassle.<br /><br />The downsides:<br />1. The "flapper" oval foam seal that reduces noise and, more importantly, seals the vent when closed came off immediately. With the first use, it went flying off. The important thing is that without this foam, there is a decent sized gap at the top of the flapper that bugs and outside air could get into. I reattached it with some silicon and hopefully that does the trick. The manufacturer needs to use a better adhesive.<br />2. The rubber seal at the base makes installation a challenge. It is installed with friction and comes off with the slightest bump. I had to make quite an effort to keep it in place during install. Note: I wish I had glued it on the day before, so it stayed in place during install rather than fighting it. The stainless steel construction is top-notch and the damper works perfectly for dryer vent use. However, the foam seal came off easily and the rubber seal at the base made installation difficult. Overall, a great product with minor flaws that can be easily fixed with some DIY solutions.
Video doorbell I’ve had a lot of frustration with this product and was considering returning it. It required a lot of troubleshooting after installation based on what was within its view. I suspect the original firmware had problems with audio/video quality that was fixed with a factory reset and registration as a new doorbell. A WiFi router sets about five feet behind the doorbell. Motion detection was triggered at sunrise every day, waking my dog and I up, and I was faced with a puppy who wanted to interact with me an hour before my alarm went off. This was fixed a few weeks later by installing a wedge kit to angle the camera so a wall was not in view.<br /><br />Pros - Easy setup. Get notified by phone. Easy to download video of events. Able to talk with people. Can setup active motion zones. Night vision. Fish eye lens covers a large area. Can snooze motion detection while you cut grass or hang out on the porch. Can set schedule of when device operates. Friendly staff via email/twitter.<br /><br />Cons - Initial problems with audio/video quality. Walls trigger motion detection at sunrise regardless of motion zones setup within the app. Limited choice of sounds on device and smartphones. I initially had issues with this product but after troubleshooting and adjusting settings, it improved. Easy setup, notifications, video download, two-way talk, motion zones, night vision, wide coverage, snooze feature, scheduling, and helpful customer service. However, initial audio/video problems and motion detection triggered by walls at sunrise were frustrating. Limited sound options.
Radio Overall it’s a well working radio. Sound is decent and clear. Display is nice looking. However the only knob is a multi function nightmare. Gone are the days of just reaching over to adjust squelch or volume with a quick turn of a knob. You got to choose the function then turn knob. Really irritating at least. The function knob is also very stiff and doesn’t turn smoothly. This radio is getting pulled out of main driver and thrown into my pickup. I really don’t like it. A shame Anytone was unable to give you two knobs, Volume and Squelch. Recommendation: Spend an extra hundred and buy a better radio.<br />Follow-up: 9/2/18 - I decided to give it a second chance. I pulled this radio off the shelf and immersed myself into programming some frequencies by hand. Took some time figuring out the quirks, like you need to program UHF frequencies when your on B radio. If you try to do it on A, your split will be 600 mhz instead of 5 khz. Well after all this learning I decided to mount it in my car. Its really not to bad once its programmed. I would however recommend a larger external speaker though since the small size radio just cant put out good sound. The other irritating thing is Channel names are limited to 5 characters. Yikes. Initially disappointed with the single function knob, but after giving it a second chance and learning its quirks, I found it to be a decent radio once programmed. Sound quality is lacking, but overall not bad for the price. Would recommend a larger external speaker for better sound.
Reusable K-Cup This isn’t compatible with some machines, out of the box. Keurig has been increasing the size of the top needle, or possibly moving it slightly off-center on some models. When using a K-Cup, that never matters of course. When using a reusable K-Cup like this one, the machine’s lid might not close without ruining the machine … or without wearing down either the needle or the reusable K-Cup’s lid opening.<br /><br />They’re all copying each other’s designs, because of course they are. Who’s gonna be brave, and the first to just make the top opening a radical 1 mm wider?<br /><br />The "smoosh down" method of securing the lid isn’t the best, however I’m aware they’re all like that. If using a disposable filter, push it down into the cup, otherwise the lid will not stay shut. Accept maybe a little bit of sediment in your coffee mug as a result. Yeah, I know, that’s what a disposable paper filter is supposed to mitigate. This reusable K-Cup may not fit all machines perfectly, but works fine with K-Cups. The design is similar to others, but could benefit from a wider top opening. The lid may need extra pressure to close, especially with a disposable filter. Overall, it gets the job done with minor inconveniences.
Headlamp I got this headlamp mainly for trail running in the dark. It’s a bright light that has more of a surround light. There are different settings for brightness. There is also the ability to turn on just the center lights which is better for focusing on what is in front of me on the trail. These are the pros on the light.<br /><br />Unfortunately, there are cons that are big enough that made me drop this light a couple of stars. First of all, it’s stiff on my head and not really comfortable for trail running or hiking. The battery pack sits kind of heavy on the back of my head. I’m not a fan at all of the charging system with the prongs that need to line up just right. It would have been so much easier to make this a regular USB charging system that would have been compatible with other charging cords.<br /><br />This will be going in my camping trailer and used for camping. On the plus side, this will be one of the brightest headlamps I’ve used for camping and will make a great addition to my camp supplies. This would also work really well for those working outside in the dark when brightness is more important than comfort. This headlamp is bright with different settings, great for trail running. However, it’s uncomfortable and has a finicky charging system. Despite the cons, it’s perfect for camping and outdoor work where brightness is key.
Headset I have the Jabra 35 headset also and this one is noticably different with regard to the quality of phone calls, but that’s really the only material difference I can find in my day to day use. The other unit does offer some noise cancellation effects that this one doesn’t have and people you call can definitely tell the difference. My husband said that he could hear background noise much more clearly when I called him from this 15 than with the 35.<br /><br />As far as battery life, this thing is really solid. I’ve used for three hours on two different days without recharging between days. I primarily use it to listen to audiobooks while I’m doing my barn chores and checking fence lines. I have noticed that I’m pretty limited with the distance I can stray from my physical phone before this headset begins to stutter to lose signal - it seems to be about 20 feet with my Sony Xperia phone. The Jabra 15 headset has great call quality but lacks noise cancellation. Battery life is impressive, lasting three hours over two days. However, signal range is limited to about 20 feet from my phone. Overall, a solid choice for calls and audiobooks, but lacking in some features compared to the 35.